Compute the average speed of water in a pipe having an i.d. of 5cm and delivering 2.5m^3 of water per hour.
1
Expert's answer
2018-04-18T08:12:07-0400
dV/dt=Av
dV/dt=2.5 m^3/h=2.5/3600 m^3/s
d=5 cm=0.05 m
The average speed of water in a pipe is
v=(dV/dt)/A=(dV/dt)/((πd^2)/4)=(2.5/3600)/(π/4 (0.05)^2 )=0.35 m/s.